# Towards Understanding Variants of Invariant Risk Minimization through the Lens of Calibration

### v1

@@ 5.3.2 Results
RMNIST ACC 91.1$\pm$1.0 91.9$\pm$1.0 90.9$\pm$0.9 89.1$\pm$0.4 90.8$\pm$1.0 87.4$\pm$0.8
ECE 5.27$\pm$1.11 3.98$\pm$0.69 1.58$\pm$0.44 6.93$\pm$0.56 7.17$\pm$0.98 7.88$\pm$0.79
PACS ACC 73.6$\pm$2.0 74.9$\pm$1.9 75.9$\pm$0.9 71.5$\pm$1.7 73.2$\pm$1.3 76.3$\pm$1.2
ECE 11.99$\pm$1.86 13.35$\pm$1.45 12.85$\pm$0.60 13.73$\pm$1.93 12.50$\pm$1.45 12.46$\pm$0.95
VLCS ACC 70.3$\pm$1.0 70.5$\pm$1.2 69.4$\pm$0.7 68.8$\pm$0.8 69.9$\pm$1.2 71.1$\pm$1.2
ECE 12.53$\pm$1.42 11.90$\pm$1.26 7.84$\pm$1.55 10.57$\pm$1.81 8.70$\pm$0.93 10.34$\pm$1.11
Avg. ACC 78.3$\pm$1.3 79.1$\pm$1.4 78.8$\pm$1.12 76.5$\pm$0.8 78.1$\pm$1.2 78.3$\pm$1.1
ECE 9.93$\pm$1.46 9.74$\pm$1.13 7.42$\pm$0.86 10.41$\pm$1.43 9.46$\pm$1.12 10.23$\pm$0.95
@@
Figure 5: The relationship between out-of-distribution accuracy and ECE varies with the $\lambda$ parameter of IRMv1 in the RMNIST dataset. As the color shifts from blue to green, $\lambda$ increases, indicating stronger regularization in the model. Points closer to blue, situated on the middle right of the figure, correspond to intermediate ECE and high accuracy. In contrast, points nearer to green, located towards the lower left, indicate better ECE but lower accuracy. This suggests that in more practical datasets, such as those with diversity shifts, stronger regularization may improve ECE, but this often comes at the cost of a significant decrease in accuracy.
The average results from three different random seeds under the aforementioned conditions are presented in Table 2, where although all methods achieved a similar average OOD accuracy, notable differences were observed in OOD ECE. Specifically, the ECE of IB-IRM was on average lower than that of any other method. Additionally, Figure 4 visualizes the relationship between ECE in the training and test environments, with the red solid line representing cases where ECE is equal in both environments. Ideally, models should be distributed near this red line from the perspective of cross-environmental calibration of uncertainty. Specifically, in Figures 3(a) and 3(c), IB-IRM is distributed near the red solid line compared to other methods, suggesting that it does not overfit to the training environment and exhibits more ideal feature learning.
@@ 6 Discussion and Conclusion
In this paper, focusing on IRM, which is challenging to realize due to the difficulties of bi-level optimization, we conducted comparative evaluations using ECE, a computationally feasible metric for calibration, to better understand its approximation methods.
The experimental results suggested that IB-IRM, which applies the information bottleneck method to IRM, is superior in terms of calibration. It is hypothesized that IB-IRM prevents excessively high confidence levels caused by dependence on environment-specific features by controlling the complexity of the feature extractor $\Phi(X)$. Additionally, it was suggested that regularization penalties introduced for the purpose of OOD generalization are effective in achieving lower OOD ECE. However, on the other hand, it was observed that in more practical datasets (Diversity Shift), while the ECE of the model becomes lower, the OOD accuracy worsens. This indicates a trade-off between OOD ECE and accuracy, suggesting that careful balancing of these two metrics is important in aiming for OOD generalization.
As a future challenge, it has become clear that the degree of invariant feature learning cannot be fully measured by accuracy in out-of-environment data alone, necessitating the establishment of broader generalization metrics and measurement methods. Furthermore, a better understanding of the relationship between OOD generalization and calibration, and the development of new approximation methods for IRM based on this understanding, are anticipated.
@@ Acknowledgments
This work used computational resources TSUBAME3.0 supercomputer provided by Tokyo Institute of Technology through the TSUBAME Encouragement Program for Younger Users awarded to Kotaro Yoshida.

### v2

@@ 5.3.2 Results
Figure 4: The relationship between OOD accuracy and ECE varies with the $\lambda$ parameter of IRMv1 in the RMNIST dataset. As the color shifts from blue to green, $\lambda$ increases, indicating stronger regularization in the model. Points closer to blue, situated on the middle right of the figure, correspond to intermediate ECE and high accuracy. In contrast, points nearer to green, located towards the lower left, indicate better ECE but lower accuracy. This suggests that in more practical datasets, such as those with diversity shifts, stronger regularization may improve ECE, but this often comes at the cost of a significant decrease in accuracy.
The average results from three different random seeds under the aforementioned conditions are presented in Table 2, where although all methods achieved a similar average OOD accuracy, notable differences were observed in OOD ECE. Specifically, the ECE of IB-IRM was, on average, lower than that of any other method. Additionally, Figure 5(c) visualizes the relationship between ECE in the training and test environments, with the red solid line representing cases where ECE is equal in both environments. Ideally, models should be distributed near this red line from the perspective of cross-environmental calibration of uncertainty. Specifically, in Figures 5(a) and 5(c), IB-IRM is distributed near the red solid line compared to other methods, suggesting that it does not overfit to the training environment and exhibits more ideal feature learning.
@@
{mdframed}
[linecolor=white, outerlinewidth=1pt, roundcorner=5pt] ERM IRMv1 BIRM IB-IRM IRM Game PAIR ID-ECE = OOD-ECE
@@
OOD ACC RMNIST 91.1
$\pm$1.0
91.9
$\pm$1.0
90.9
$\pm$0.9
89.1
$\pm$0.4
90.8
$\pm$1.0
87.4
$\pm$0.8
PACS 73.6
$\pm$2.0
74.9
$\pm$1.9
75.9
$\pm$0.9
71.5
$\pm$1.7
73.2
$\pm$1.3
76.3
$\pm$1.2
VLCS 70.3
$\pm$1.0
70.5
$\pm$1.2
69.4
$\pm$0.7
68.8
$\pm$0.8
69.9
$\pm$1.2
71.1
$\pm$1.2
Avg. 78.3
$\pm$1.3
79.1
$\pm$1.4
78.8
$\pm$1.12
76.5
$\pm$0.8
78.1
$\pm$1.2
78.3
$\pm$1.1
OOD ECE RMNIST 5.27
$\pm$1.11
3.98
$\pm$0.69
1.58 $\pm$0.44 6.93
$\pm$0.56
7.17
$\pm$0.98
7.88
$\pm$0.79
PACS 11.99 $\pm$1.86 13.35
$\pm$1.45
12.85
$\pm$0.60
13.73
$\pm$1.93
12.50
$\pm$1.45
12.46
$\pm$0.95
VLCS 12.53
$\pm$1.42
11.90
$\pm$1.26
7.84 $\pm$1.55 10.57
$\pm$1.81
8.70
$\pm$0.93
10.34
$\pm$1.11
Avg. 9.93
$\pm$1.46
9.74
$\pm$1.13
7.42 $\pm$0.86 10.41
$\pm$1.43
9.46
$\pm$1.12
10.23
$\pm$0.95
@@
6 Information Bottleneck in Calibration
In our experiments, we observe that IRM with information bottleneck is effective in calibration, as shown in Figure 2 and Figure 5(c). Therefore, in this section, we provide the results of the ablation study for IBIRM.
First, we investigated the impact of the penalty term related to the information bottleneck in IBIRM on calibration. Figure 6 visualizes various evaluation metrics in CMNIST when altering the strength of the information bottleneck in IBIRM, that is, the magnitude of $\gamma$ in (7). As shown in panel (a), which visualizes this based on OOD ECE and OOD Accuracy, we confirm the clear trend that strengthening the regularization of $\gamma$ improves both accuracy and ECE. Panel (b) shows the relationship between ECE in ID and OOD. Increasing $\gamma$ aligns both ECEs towards the red line, representing equality for conditional probability. In other words, this constitutes a successful calibration across environments. These findings suggest that not only is the IRMv1’s regularization term aimed at OOD scenarios, but the information bottleneck also further promotes the learning of invariant features.
Additionally, we demonstrate the validity of our results based on the theoretical principles of the information bottleneck. In OOD scenarios, the objective of the information bottleneck is to minimize the mutual information $I(X;\Phi(X))$ between the input $X$ and the intermediate representation $\Phi(X)$ while preserving the mutual information $I(\Phi(X);Y)$ between $\Phi(X)$ and the output $Y$. Intuitively, this means learning $\Phi(X)$ to reduce as much information from $X$ as possible while retaining as much information about $Y$. When $\Phi(X)$ is a deterministic transformation of $X$, the entropy $H(\Phi(X))$ can be used to minimize $I(X;\Phi(X))$ (Kirsch et al., 2021). In the context of IRM, which aims to learn invariant features, keeping $H(\Phi(X))$ small encourages $\Phi$ to capture invariant features $X_{inv}$ from $X$ and discard spurious features $X_{sup}$. In fact, Ahuja et al. (2022) theoretically demonstrated, using realistic SEMs, that minimizing $H(\Phi(X))$ in IRM leads to $\Phi$ discarding spurious features and grasping invariant features. Our results empirically support the effectiveness of applying the information bottleneck method to invariant feature learning.
{mdframed}
[linecolor=white, outerlinewidth=1pt, roundcorner=5pt] ID-ECE = OOD-ECE
Figure 6: The figures illustrate the impact of the information bottleneck on calibration within IBIRM. Panel (a) displays the relationship between ECE and Accuracy in the OOD context, investigated by varying the coefficient $\gamma$ of the information bottleneck penalty in the IBIRM formulation (7). As the value of $\gamma$ increases, both metrics show improvement. Panel (b) visualizes the values of ECE in both ID and OOD. Same as (a), altering the value of $\gamma$ shows that the larger the value, the more the data aligns with the red line, which indicates equality between the two ECEs. With a sufficiently large $\gamma$, the points are almost perfectly distributed along the red line, indicating successful calibration across multiple environments.
@@ 7 Discussion and Conclusion
In this paper, we focus on IRM, which is challenging to realize due to the difficulties of bi-level optimization. We conducted comparative evaluations using ECE, a computationally feasible metric for calibration, to understand its approximation methods better.
The experimental results suggested that IB-IRM, which applies the information bottleneck method to IRM, is superior in calibration. Additionally, it was suggested that regularization penalties introduced for OOD generalization effectively achieve lower OOD ECE. However, on the other hand, it was observed that in more practical datasets (Diversity Shift), while the ECE of the model becomes lower, the OOD accuracy worsens. This indicates a trade-off between OOD ECE and accuracy, suggesting that carefully balancing these two metrics is important in aiming for OOD generalization.
As a future challenge, it has become clear that the degree of invariant feature learning cannot be fully measured by accuracy on OOD data alone, necessitating the establishment of systematic generalization metrics and measurement methods. Furthermore, a better understanding of the relationship between OOD generalization and uncertainty calibration, as well as the development of novel approximation methods for IRM based on this understanding, is anticipated. In particular, gaining a deeper understanding of the impact of the information bottleneck on OOD generalization and calibration will be key to addressing the issues of OOD generalization and calibration. A more comprehensive grasp of how the information bottleneck principle affects the model’s ability to generalize to unseen environments and maintain well-calibrated predictive uncertainties will be crucial in developing more robust and reliable models.
@@ Acknowledgments
We acknowledge the generous allocation of computational resources from the TSUBAME3.0 supercomputer facilitated by the Tokyo Institute of Technology. This assistance came through the TSUBAME Encouragement Program for Young / Female Users, whose support was instrumental to this research.
Appendix
Appendix A Details of experiments
A.1 Implementation of IRM Variants
This section provides a detailed description of each method used in the experiments.
A.1.1 IRMv1
The learning process was conducted using the loss calculated based on the formulation shown in 6. The hyperparameter $\lambda$ is used to adjust the impact of the IRMv1 penalty.
A.1.2 IBIRM
The learning process was performed using the loss calculated based on the formulation shown in 7. The hyperparameters $\lambda$ and $\gamma$ are used to adjust the influence of the IRMv1 penalty and the information bottleneck penalty, respectively.
A.1.3 IRM Game
IRM Game has two variants: F-IRM Game and V-IRM Game. The difference between them lies in whether the featurizer $\Phi$ is fixed or not. F-IRM Game fixes $\Phi=I$ as an identity matrix, while V-IRM Game considers a variable Φ. For each dataset, the variant that achieved higher performance in terms of accuracy was selected.
A.1.4 PAIR
PAIR employs a weighted average of ERM loss, IRMv1 loss, and VREx loss as the final loss for learning. When calculating the weights, a hyperparameter called $preference$ is used to adjust the scaling of each loss.
A.1.5 BIRM
In BIRM, when calculating the IRMv1 loss, Monte Carlo sampling is used to sample $N$ times from the posterior distribution of the parameters, and the average of these samples is used to compute the loss. This approach takes into account the uncertainty of the parameters when calculating the loss. The standard deviation of the Gaussian noise added to the parameters during sampling is treated as a hyperparameter denoted as $birm\_sd$. In all experiments, $N$ was set to 5.
A.2 Datasets
A.2.1 Domainbed
We conducted experiments using Colored MNIST (CMNIST) (Arjovsky et al., 2020), Rotated MNIST (RMNIST) (Ghifary et al., 2015), PACS (Li et al., 2017), and VLCS (Fang et al., 2013) datasets from DomainBed (Gulrajani & Lopez-Paz, 2020). We split each dataset into training and validation sets, with 80% used for training and the remaining 20% for validation. The environment partitions were as follows:
  • CMNIST: $E_{train}=[10\%,20\%]$, $E_{test}=[90\%]$
  • RMNIST: $E_{train}=[15^{\circ},30^{\circ},45^{\circ},60^{\circ},75^{\circ}]$, $E_{test}=[0^{\circ}]$
  • PACS: $E_{train}=[Photo,Painting,Sketch]$, $E_{test}=[Art]$
  • VLCS: $E_{train}=[Caltech101,LabelMe,SUN09]$, $E_{test}=[VOC2007]$
A.2.2 Hyperparameters
In the experiments, we set the batch size to 256 for CMNIST, 128 for RMNIST, and 16 for PACS and VLCS. Grid search was performed on the learning rate for all experiments, with values of [1e-4, 5e-4, 1e-3, 5e-3]. For the hyperparameters specific to each approximation method, grid search was conducted as shown in Table 3.
Parameter
$\lambda$ [1, 1e1, 1e2, 1e3, 1e4]
$\gamma$ [1, 1e1, 1e2, 1e3, 1e4]
$preference$ [0, 1, 2, 3, 4]
$birm\_sd$ [5e-2, 1e-1, 2e-1]
Table 3: Hyperparameters for each method
Model selection was performed using an oracle based on accuracy in the test environment for Correlation shift, while for Diversity shift, it was based on validation accuracy in the training environment. Correlation shift is known to be a relatively challenging task(Ye et al., 2022) as it involves a shift in the conditional probability $P[Y|X]$, where $X$ represents the input data and $Y$ represents the corresponding ground truth labels. As shown in Figure 1(b), there exists a trade-off between accuracy in the training and test environments for all methods, necessitating oracle model selection. In contrast, for Diversity shift, there is a correlation between ID and OOD accuracy across all methods as shown in Figure 9(c), allowing for model selection based on performance in the training environment.
Appendix B Additional results
B.1 Impact of penalties of IBIRM
Figure 3 in CMNIST compares the OOD ECE of the model when varying the regularization penalty $\lambda$ in IRMv1. Additionally, it compares the OOD ECE of the model when varying the coefficients of the two regularization penalties in IBIRM 7. Similar to 3, as the regularization penalty increases, the OOD ECE decreases.
(a) OOD Accuracy
(b) OOD ECE
$\gamma$ in IBIRM
1 10 100 1000 10000
Figure 7: Comparison of the model’s OOD ECE in CMNIST, with various values of $\gamma$ and $\lambda$ fixed at a sufficiently large value (10000) in the formulation of IBIRM 7. It was observed that as $\gamma$ increases, i.e., as the regularization penalty from the information bottleneck method becomes more significant, the ECE of the model in OOD scenarios decreases.
(a) OOD Accuracy
(b) OOD ECE
$\lambda$ in IBIRM
1 10 100 1000 10000
Figure 8: Comparison of the model’s OOD ECE in CMNIST, with various values of $\lambda$ and $\gamma$ fixed at a sufficiently large value (10000) in the formulation of IBIRM 7. It was observed that as $\lambda$ increases, i.e., as the regularization penalty from IRMv1 in IB-IRM becomes more significant, the ECE of the model in OOD decreases.
B.2 Relationship between ID and OOD accuracy in diversity shift
As shown in Figure 1(b), in the presence of correlation shift, including the CMNIST dataset, a trade-off exists between ID and OOD accuracy for all methods. However, as illustrated in Figure 9(c), for diversity shift, a positive correlation between ID and OOD accuracy was observed across all methods.
(a) RMNIST
(b) PACS
(c) VLCS
Figure 9: This figures illustrate the relationship between ID and OOD accuracy on each dataset of Diversity shift. It is evident that there exists a positive correlation between accuracy on the ID data and accuracy on the OOD data.
{mdframed}
[linecolor=white, outerlinewidth=1pt, roundcorner=3pt, innerleftmargin=0pt, innerrightmargin=0pt] ERM IRMv1 BIRM IB-IRM IRM Game PAIR
Appendix C Additional discussion and potential future directions
The experimental results demonstrated that IBIRM, which employs an information bottleneck, excels in terms of model calibration. This can be attributed to the fact that applying the information bottleneck restricts the representational power of the featurizer $\Phi(X)$, limiting its complexity and enabling it to discard spurious features, thereby reducing dependence on them. This approach of constraining the complexity of data representations has been applied in various contexts, and there are related studies in the field of Large Language Models (LLMs) as well.
Ruan et al. (2021) demonstrated that applying the information bottleneck method to the vision-language model CLIP (Radford et al., 2021) improved its OOD generalization performance. Furthermore, Hu et al. (2021) proposed a new fine-tuning method called Low-Rank Adaptation (LoRA). This method involves attaching a very low-rank linear layer as an adapter to a pre-trained model and fine-tuning only that part for a specific task, achieving performance comparable to full fine-tuning with significantly lower memory requirements. This approach can be interpreted as compressing task-specific information into a low-rank representation during learning, suggesting a connection with the information bottleneck method.
Based on the aforementioned research, future studies are anticipated to investigate the impact of learning through information compression on OOD generalization and calibration, as well as to develop novel methods that simultaneously address these two issues.